OpenStudy (anonymous):

radicals: √3 * √6 i got 3√2, is this right?

jimthompson5910 (jim_thompson5910):

You are correct, \[\sqrt{3}\sqrt{6}=\sqrt{18}=\sqrt{9*2}=\sqrt{9}\sqrt{2}=3\sqrt{2}\] So \[\sqrt{3}\sqrt{6}=3\sqrt{2}\]

OpenStudy (anonymous):

thank youuuu!